# Positive Semidefinite Supermartingales and Randomized Matrix Concentration Inequalities

### v3

@@ 4.4 Matrix-Valued Submartingales and Backward Supermartingales
Let $\{\mathbf{Y}_{n}\}$ be an $\mathcal{S}_{d}$-valued submartingale, and $\tau$ be a stopping time on $\{\mathcal{F}_{n}\}$. Then, $\{\mathbf{Y}_{n\wedge\tau}\}$ is also a submartingale on $\{\mathcal{F}_{n}\}$. Consequently, if $\tau\leqslant N$ almost surely for some $N\in\mathbb{N}$, then $\mathbb{E}\mathbf{Y}_{\tau}\preceq\mathbb{E}\mathbf{Y}_{N}$.
@@
Next, we illustrate how Theorem 4.14 remains valid even without assuming that the process $\{\mathbf{Y}_{n}\}$ starts at $n=0$ — or starts at all; it is applicable even if $\{\mathbf{Y}_{n}\}$ is indexed on $n\in\{\dots,-2,-1,0,1,\dots,N\}$. We formalize this by considering a backward filtration $\{\mathcal{G}_{n}\}$, i.e., $\mathcal{G}_{0}\supseteq\mathcal{G}_{1}\supseteq\dots$, and an adapted backward submartingale $\{\mathbf{Y}_{n}\}$, i.e., $\mathbf{Y}_{n}$ is $\mathcal{G}_{n}$-measurable and $\mathbb{E}[\mathbf{Y}_{n}|\mathcal{G}_{n+1}]\succeq\mathbf{Y}_{n+1}$, while keeping the time indices $n\geqslant 0$. This, in the scalar regime, sets the stage for the backward Ville’s inequality which has been shown equivalent to the (forward) Doob’s inequality (hence the duality between Ville’s and Doob’s inequality; see the second proof of Theorem 2 in Manole and Ramdas [2023]), for which we have the following matrix version.
@@
Let $\{\mathbf{Y}_{n}\}$ be an $\mathcal{S}_{d}^{+}$-valued backward submartingale adapted to the backward filtration $\{\mathcal{G}_{n}\}$. For any $\mathbf{A}\in\mathcal{S}_{d}^{++}$,
@@
Take an arbitrary $N\in\mathbb{N}$ and consider, indexed by $0\leqslant n\leqslant N$, the process $\mathbf{P}_{n}=\mathbf{Y}_{N-n}$ on the (forward) filtration $\mathcal{F}_{n}=\mathcal{G}_{N-n}$. It is clear that $\{\mathbf{P}_{n}\}$ is a (forward) submartingale on $\{\mathcal{F}_{n}\}$, to which we can apply Theorem 4.14,
@@
$\{\exists n,\;\mathbf{Y}_{n}\npreceq\mathbf{A}\}=\liminf_{N\to\infty}\left\{\exists n\leqslant N,\;\mathbf{Y}_{n}\npreceq\mathbf{A}\right\}=\bigcup_{k=1}^{\infty}\bigcap_{N=k}^{\infty}\left\{\exists n\leqslant N,\;\mathbf{Y}_{n}\npreceq\mathbf{A}\right\}.$ (52)
4.5 Exchangeable Matrices
We now have the tools to prove a matrix version of the exchangeable Chebyshev inequality (XCI) we mentioned earlier. Recall from Fact 2 that when $p\in[1,2]$, the function $\mathbf{X}\mapsto\mathbf{X}^{p}$ is operator convex within the domain $\mathcal{S}_{d}^{+}$ (for $21\leqslant p<$ 2) or $\mathcal{S}_{d}$ (for $p=2$) respectively,
@@ 4.5 Exchangeable Matrices
Applying the matrix backward Ville’s inequality (Theorem 4.15) to this process $\{\mathbf{R}_{n}\}$, we immediately have the following two matrix extensions of (XCI).
@@
$\mathbb{P}(\exists n,\;\operatorname{abs}(\overline{\mathbf{X}}_{n}-\mathbf{M})\npreceq\mathbf{A})\leqslant\mathrm{tr}(\mathbf{V}\mathbf{A}^{-2}).$ (XMCI)
@@
Now use the operator monotonicity of $\mathbf{X}\mapsto\mathbf{X}^{1/2}$,
@@
concluding the proof of (XMCI). To show (58), simply apply Theorem 4.15 to the submartingale $\mathbf{R}_{n}=(\overline{\mathbf{X}_{n}}-\mathbf{M})^{2}$ with $n\geqslant N$, obtaining
@@
due to Theorem 3.5. ∎
@@
Applying Theorem 4.15 to $\{\mathbf{R}_{n}\}$, with $\mathbf{C}$ being 00, in Lemma 4.16,
@@
Now use the operator monotonicity of $\mathbf{X}\mapsto\mathbf{X}^{1/p}$,
@@
Indeed, we can weaken the second-moment assumption to a $p$th moment, leading to, however, a concentration not around the mean but a Markov-type upper bound. This is largely because the scalar map $x\mapsto|x|^{p}$ is convex but not its matrix extension, and so the raw $p$th moment is used above. Nonetheless, due to Fact 4, the following trace inequality holds.
Therefore,
is a real-valued backward submartingale. With the scalar backward Ville’s inequality, this leads to the following anytime-valid extension of Lemma 3.7.
@@
Let $\{\mathbf{X}_{n}\}$ be an exchangeable sequence of $\mathcal{S}_{d}$-valued random matrices with mean $\mathbf{M}$ and central $p$th moment $\mathbf{V}_{p}=\mathbb{E}\operatorname{abs}(\mathbf{X}_{1}-\mathbf{M})^{p}$, where $p\geqslant 1$. Then, for any $a>0$,
@@
$\mathbb{P}(\exists n,\;\|\overline{\mathbf{X}}_{n}-\mathbf{M}\|_{\mathrm{op}}\geqslant a)\leqslant\mathbb{P}\left(\exists n,\;\mathrm{tr}\left(\operatorname{abs}(\overline{\mathbf{X}}_{n}-\mathbf{M})^{p}\right)\geqslant a^{p}\right)\leqslant a^{-p}\mathrm{tr}\mathbf{V}_{p}.$ (67)
@@
The bound above differs from Theorem 4.18 in that the concentration is now around the mean $\mathbf{M}$.
5 Randomized Matrix Bounds via Scalar e-Processes
5.1 A Conversion Lemma from Matrix to Scalar Supermartingales
We have studied the concentration of random matrices via matrix e-statistics, i.e. random matrices with bounded expectations, matrix supermartingales, matrix e-values, and matrix backward submartingales, since these quantities yield concentration inequalities via our Markov and Ville’s inequalities for matrices, oftentimes in randomized forms. Some further randomized matrix concentration inequalities can be obtained by revisiting some bounds obtained via scalar e-quantities that already exist in the literature.
@@ 5.1 A Conversion Lemma from Matrix to Scalar Supermartingales
As we mentioned earlier in Section 4.2, Howard et al. [2020, Section 6.5] prove, under different conditions, numerous inequalities for symmetric matrix martingale difference $\{\mathbf{Z}_{n}\}$ in the style of
@@
At this step, our previous methods and that of Howard et al. [2020] had diverged. We have been able to directly construct matrix-valued supermartingales in Section 4.2, leading potentially to computable sequential tests by virtue of our new (UMVI). Howard et al. [2020], on the other hand, construct scalar e-processes from the condition (68). Let us rephrase their “master lemma” behind this construction.
@@
The proof of this lemma can be found in Section B.5. Note that the upper bounded expression (70) is an instance of e-processes, introduced earlier in Section 4.3. We shall see soon that both (69) and (70) can be useful.
5.2 Examples: Self-Normalization, Bernstein, and Hoeffding
Recall from Example 4.8 the self-normalized bound, that if $\{\mathbf{X}_{n}\}$ are symmetric matrices adapted to $\{\mathcal{F}_{n}\}$ with conditional means $\{\mathbf{M}_{n}\}$ and predictable conditional variance upper bounds $\{\mathbf{V}_{n}\}$,
$\mathbb{E}\left(\exp(\gamma_{n}(\mathbf{X}_{n}-\mathbf{M}_{n})-\gamma_{n}^{2}(\mathbf{X}_{n}-\mathbf{M}_{n})^{2}/6)\mid\mathcal{F}_{n-1}\right)\preceq\exp\left(\frac{\gamma_{n}^{2}}{3}\mathbf{V}_{n}\right),$ (71)
an instance of (68). By Lemma 5.1, this implies the following statement whose weaker version is stated as Lemma 3(f) by Howard et al. [2020].
Example 5.2 (Matrix Self-Normalized Scalar Supermartingale).
Under the same assumptions as Example 4.8, the process
$L_{n}^{\textsf{SN}}=\mathrm{tr}\exp\left(\sum_{i=1}^{n}\gamma_{i}(\mathbf{X}_{i}-\mathbf{M}_{i})-\sum_{i=1}^{n}\frac{\gamma_{i}^{2}}{6}((\mathbf{X}_{i}-\mathbf{M}_{i})^{2}+2\mathbf{V}_{i})\right)$ (72)
is a supermartingale.
The proof is simply taking $\mathbf{Z}_{n}=\gamma_{n}(\mathbf{X}_{n}-\mathbf{M}_{n})$, $\mathbf{C}_{n}=\frac{\gamma_{n}^{2}}{6}(\mathbf{X}_{n}-\mathbf{M}_{n})^{2}$, and $\mathbf{C}_{n}^{\prime}=\frac{\gamma_{n}^{2}}{3}\mathbf{V}_{n}$ in Lemma 5.1. Some remarks on this result are in order. First, the expression of the scalar supermartingale $\{L_{n}^{\textsf{SN}}\}$ explains the name “self-normalized”, as an empirical variance term exists within the bound. Second, unlike the scalar case where a closed-form confidence sequence can be solved from the supermartingale combining (VI) if the means are a constant [Wang and Ramdas, 2023a, Section 5], in our matrix case, we only have an implicit bound when applying (UVI).
Corollary 5.3 (Uniformly Randomized Stopped Matrix Self-Normalized Bound).

### v4

@@ 4.1 Submartingales and Maximal Inequalities: From Forward to Backward
Let $\{\mathbf{Y}_{n}\}_{n\geqslant 0}$ be an $\mathcal{S}_{d}$-valued submartingale, and $\tau$ be a stopping time on $\{\mathcal{F}_{n}\}_{n\geqslant 0}$. Then, $\{\mathbf{Y}_{n\wedge\tau}\}_{n\geqslant 0}$ is also a submartingale on $\{\mathcal{F}_{n}\}_{n\geqslant 0}$. Consequently, if $\tau\leqslant N$ almost surely for some $N\in\mathbb{N}$, then $\mathbb{E}\mathbf{Y}_{\tau}\preceq\mathbb{E}\mathbf{Y}_{N}$.
@@
The backward Ville’s inequality comes from the idea that Doob’s inequality remains valid even without assuming that time starts at $n=0$ — or that it starts at all. In the scalar regime, this idea proves the backward Ville’s inequality (7) and proves that it is equivalent to the (forward) Doob’s inequality (hence the “duality” between Ville’s and Doob’s inequality; see the second proof of Theorem 2 in Manole and Ramdas [2023]).
This idea is valid for matrices as well; that is, Theorem 4.2 is applicable even if $\{\mathbf{Y}_{n}\}$ and $\{\mathcal{F}_{n}\}$ are indexed on $n\in\{\dots,-2,-1,0,1,\dots,N\}$. We formalize this by recalling that a backward filtration $\{\mathcal{G}_{n}\}$ is one that satisfies $\mathcal{G}_{0}\supseteq\mathcal{G}_{1}\supseteq\dots$ and an adapted backward submartingale $\{\mathbf{Y}_{n}\}$ is one such that each $\mathbf{Y}_{n}$ is $\mathcal{G}_{n}$-measurable and $\mathbb{E}[\mathbf{Y}_{n}|\mathcal{G}_{n+1}]\succeq\mathbf{Y}_{n+1}$, while we formally keep the time indices $n\geqslant 0$.
@@
Let $\{\mathbf{Y}_{n}\}_{n\geqslant 0}$ be an $\mathcal{S}_{d}^{+}$-valued backward submartingale adapted to the backward filtration $\{\mathcal{G}_{n}\}_{n\geqslant 0}$. For any $\mathbf{A}\in\mathcal{S}_{d}^{++}$,
@@
Take an arbitrary $N\in\mathbb{N}$ and consider, indexed by $0\leqslant n\leqslant N$, the process $\mathbf{P}_{n}=\mathbf{Y}_{N-n}$ on the (forward) filtration $\mathcal{F}_{n}=\mathcal{G}_{N-n}$. It is clear that $\{\mathbf{P}_{n}\}_{0\leqslant n\leqslant N}$ is a (forward) submartingale on $\{\mathcal{F}_{n}\}_{0\leqslant n\leqslant N}$, to which we can apply Theorem 4.2,
@@
$\{\exists n,\;\mathbf{Y}_{n}\npreceq\mathbf{A}\}=\liminf_{N\to\infty}\left\{\exists n\leqslant N,\;\mathbf{Y}_{n}\npreceq\mathbf{A}\right\}=\bigcup_{k=1}^{\infty}\bigcap_{N=k}^{\infty}\left\{\exists n\leqslant N,\;\mathbf{Y}_{n}\npreceq\mathbf{A}\right\}.\qed$ (48)
4.2 Application: Concentration of Exchangeable Matrices
An immediate application of Theorem 4.3 is numerous matrix Chebyshev-type concentration bounds for exchangeable random matrices. We recall that a sequence of random variables $X_{1},X_{2},\dots$ taking values in a Polish space is exchangeable if the joint distribution of $(X_{i_{1}},\dots,X_{i_{k}})$ equals that of $(X_{j_{1}},\dots,X_{j_{k}})$ for any distinct $i_{1},\dots,i_{k}$ and distinct $j_{1},\dots,j_{k}$. Repetitions of a single random variable, the Pólya urn model, sampling from a finite set without replacement, as well as i.i.d. random variables are all examples of exchangeable sequences.
The scalar exchangeable Chebyshev inequality was recently proved by Ramdas and Manole [2024, Theorem 2]. Recall from Fact 1 that when $p\in[1,2]$, the function $\mathbf{X}\mapsto\mathbf{X}^{p}$ is operator convex within the domain $\mathcal{S}_{d}^{+}$ (for $21\leqslant p<$ 2) or $\mathcal{S}_{d}$ (for $p=2$) respectively,
@@ 4.2 Application: Concentration of Exchangeable Matrices
Applying the matrix backward Ville’s inequality (Theorem 4.3) to this process $\{\mathbf{R}_{n}\}$, we immediately have the following two exchangeable matrix Chebyshev inequalities. First, the following holds when a finite variance exists. Recall that the anticommutator between two sqaure matrices $\mathbf{A},\mathbf{B}$ is defined as $\{\mathbf{A},\mathbf{B}\}=\mathbf{A}\mathbf{B}+\mathbf{B}\mathbf{A}$.
@@
$\mathbb{P}(\exists n,\;\operatorname{abs}(\overline{\mathbf{X}}_{n}-\mathbf{M})\npreceq\mathbf{A})\leqslant\mathrm{tr}(\mathbf{V}\mathbf{A}^{-2}).$ (54)
@@
Now we use the operator monotonicity of $\mathbf{X}\mapsto\mathbf{X}^{1/2}$ (Fact 1),
@@
concluding the proof of (54). To show (55), simply apply Theorem 4.3 to the submartingale $\mathbf{R}_{n}=(\overline{\mathbf{X}_{n}}-\mathbf{M})^{2}$ with $n\geqslant N$, obtaining
@@
where the last step is due to Lemma B.2. ∎
We remark that the assumption $\mathbb{E}\{\mathbf{X}_{i}-\mathbf{M},\mathbf{X}_{j}-\mathbf{M}\}=0$ for $i\neq j$ is satisfied when $\mathbf{X}_{1},\mathbf{X}_{2},\dots$ are pairwise uncorrelated, which is in turn satisfied under martingale difference assumption (i.e., $\{\mathbf{X}_{i}\}$ is adapted to some filtration $\{\mathcal{F}_{i}\}$ with $\mathbb{E}(\mathbf{X}_{i}|\mathcal{F}_{i-1})=\mathbf{M}$ and $\mathbf{Var}(\mathbf{X}_{i}|\mathcal{F}_{i-1})\preceq\mathbf{V}$), which is in turn implied if they are i.i.d.
The inequality (54) generalizes the matrix Chebyshev inequality by Ahlswede and Winter [2002, Theorem 14] which states that $\mathbb{P}(\operatorname{abs}(\overline{\mathbf{X}}-\mathbf{M})\npreceq\mathbf{A})\leqslant\mathrm{tr}(\mathbf{V}\mathbf{A}^{-2})$ for one square-integrable matrix $\mathbf{X}$ with variance $\mathbf{V}$. It is only by assuming $\mathbb{E}\{\mathbf{X}_{i}-\mathbf{M},\mathbf{X}_{j}-\mathbf{M}\}=0$ that we may obtain a “shrinking” factor $N^{-1}$ on the tail probability that shows up in the right hand side of (55). Without this assumption, these exchangeable matrices can all be the repetition of a single random matrix $\mathbf{X}_{1}=\mathbf{X}_{2}=\dots$ in which case (54) is reduced to the matrix Chebyshev inequality by Ahlswede and Winter [2002, Theorem 14]. On the other hand, with the stated additional assumption $\mathbb{E}\{\mathbf{X}_{i}-\mathbf{M},\mathbf{X}_{j}-\mathbf{M}\}=0$, the inequality (55) generalizes the matrix “weak law of large numbers” by Ahlswede and Winter [2002, Corollary 16] which we already quoted as (10), that $\mathbb{P}(\operatorname{abs}(\overline{\mathbf{X}}_{N}-\mathbf{M})\npreceq\mathbf{A})\leqslant N^{-1}\mathrm{tr}(\mathbf{V}\mathbf{A}^{-2})$ for $N$ i.i.d. matrices $X_{1},\dots,X_{N}$. Both generalizations, we note, offer the exact same concentration at the sample sizes (1 and $N$ respectively) compared to the original results by Ahlswede and Winter [2002], while also extending infinitely into all future sample sizes. These generalizations generalize those in the scalar case by Ramdas and Manole [2024, Section 2.2].
We next present results for exchangeable matrices with heavier tails.
@@
We define $\mathbf{R}_{n}=(\overline{\mathbf{X}}_{n})^{p}$, which, since $p\in[1,2]$, is a matrix backward submartingale due to Lemma 4.4. Applying Theorem 4.3 to $\{\mathbf{R}_{n}\}$, we see that
@@
Now use the operator monotonicity of $\mathbf{X}\mapsto\mathbf{X}^{1/p}$ (Fact 1),
@@
We remark that, from Theorem 4.5 to Theorem 4.6, we weaken the second moment to a $p$th moment assumption, but in an imperfect manner as we obtain a concentration not around the mean $\mathbf{M}$ but a Markov-type “upper tail bound” for $\mathcal{S}_{d}^{+}$-valued matrices instead. This is largely because the scalar map $x\mapsto|x|^{p}$ is convex on (,)($\infty$,$\infty$)( - ∞ , ∞ ) but not its matrix extension, and so the raw $p$th moment is used above. Also, Theorem 4.6 fails if $p>2$ because the scalar map $x\mapsto x^{p}$ is convex on $[0,\infty)$ but not its matrix extension. Nonetheless, thanks to Fact 3, we can resort to trace scalarization to obtain the following spectral and trace inequalities that hold for all $p\geqslant 1$.
@@
Let $\{\mathbf{X}_{n}\}$ be an exchangeable sequence of $\mathcal{S}_{d}$-valued random matrices with mean $\mathbf{M}$ and central $p$th moment $\mathbf{V}_{p}=\mathbb{E}(\operatorname{abs}(\mathbf{X}_{1}-\mathbf{M}))^{p}$, where $p\geqslant 1$. Then, for any $a>0$,
@@
$\mathbb{P}(\exists n,\;\|\overline{\mathbf{X}}_{n}-\mathbf{M}\|\geqslant a)\leqslant\mathbb{P}\left(\exists n,\;\mathrm{tr}\left(\operatorname{abs}(\overline{\mathbf{X}}_{n}-\mathbf{M})^{p}\right)\geqslant a^{p}\right)\leqslant a^{-p}\mathrm{tr}\mathbf{V}_{p}.$ (62)
@@
The inequality above, which we defer its straightforward proof to Section B.3, differs from Theorem 4.6 in that the concentration is now around the mean $\mathbf{M}$.
5 Remarks and Extensions
5.1 Tightness of Matrix Markov’s and Ville’s Inequalities
[…]
$\mathcal{C}_{\mathrm{tr}}(\alpha)=\{\mathbf{X}\in\mathcal{S}_{d}^{+}:\mathrm{tr}\mathbf{X}\leqslant d/\alpha\};$ (63)
by $\mathcal{C}_{\mathbf{v}}(\alpha)$ the concentration region by applying (MI) to $\mathbf{v}^{\mathsf{T}}\mathbf{X}\mathbf{v}$ where $\|\mathbf{v}\|=1$:
$\mathcal{C}_{\mathbf{v}}(\alpha)=\{\mathbf{X}\in\mathcal{S}_{d}^{+}:\mathbf{v}^{\mathsf{T}}\mathbf{X}\mathbf{v}\leqslant 1/\alpha\}.$ (64)
Let $\mathcal{A}_{\alpha}$ be the following set of matrices:
$\mathcal{A}_{\alpha}=\{\mathbf{A}\in\mathcal{S}_{d}^{++}:\mathrm{tr}(\mathbf{A}^{-1})=\alpha\},$ (65)
and we denote by $\mathcal{C}(\mathbf{A})$ the concentration region by applying (MMI) to $\mathbf{X}$ with threshold $\mathbf{A}\in\mathcal{A}_{\alpha}$:
$\mathcal{C}({\mathbf{A}})=\{\mathbf{X}\in\mathcal{S}_{d}^{+}:\mathbf{X}\preceq\mathbf{A}\}.$ (66)
Following e.g. Minsker [2017], we denote by $\operatorname{r}(\mathbf{A})$ the effective rank $\mathrm{tr}\mathbf{A}/\|\mathbf{A}\|\in[1,d]$ of a PD matrix. We can establish the following relations among these concentration regions via elementary linear algebra.
Proposition 5.1.
Suppose $\alpha\in(0,1)$ and $d\geqslant 2$.
  1. 1.
    $\mathcal{C}_{\mathrm{tr}}(\alpha)\subsetneq\mathcal{C}((d/\alpha)\mathbf{I})$;
  2. 2.
    If $\mathbf{A}\in\mathcal{A}_{\alpha}$ and $\mathbf{A}\neq(d/\alpha)\mathbf{I}$, then $\mathcal{C}_{\mathrm{tr}}(\alpha)\nsubseteq\mathcal{C}(\mathbf{A})$ and $\mathcal{C}(\mathbf{A})\nsubseteq\mathcal{C}_{\mathrm{tr}}(\alpha)$;
  3. 3.
    If $\mathbf{A}\in\mathcal{A}_{\alpha}$, then $\mathcal{C}_{\mathrm{tr}}(\alpha^{\prime})\subseteq\mathcal{C}(\mathbf{A})$ if and only if $\alpha^{\prime}\geqslant\frac{d}{\operatorname{r}(\mathbf{A}^{-1})}\cdot\alpha$;
  4. 4.
    If $\mathbf{A}\in\mathcal{A}_{\alpha}$, then $\mathcal{C}(\mathbf{A})\subseteq\mathcal{C}_{\mathrm{tr}}(\alpha^{\prime})$ if and only if $\alpha^{\prime}\leqslant d(\mathrm{tr}\mathbf{A})^{-1}\in[\frac{\alpha}{d},\frac{\alpha}{\operatorname{r}(\mathbf{A}^{-1})}]$;
  5. 5.
    For any $\mathbf{v}$ such that $\|\mathbf{v}\|=1$ and any $\alpha^{\prime}>\alpha$, there always exists an $\mathbf{A}\in\mathcal{A}_{\alpha^{\prime}}$ such that $\mathcal{C}(\mathbf{A})\subsetneq\mathcal{C}_{\mathbf{v}}(\alpha)$;
  6. 6.
    For any $\mathbf{A}\in\mathcal{A}_{\alpha}$ and any $\alpha^{\prime}\in(0,1)$, there exists no $\mathbf{v}$ such that $\|\mathbf{v}\|=1$ and $\mathcal{C}_{\mathbf{v}}(\alpha^{\prime})\subseteq\mathcal{C}(\mathbf{A})$.
Let us briefly unpack these statements above. The first two state that (MMI) and trace-(MI) are mutually incomparable unless one takes the threshold matrix $\mathbf{A}$ to be a multiple of $\mathbf{I}$, in which case (MMI) is weaker than trace-(MI). Statements 3 and 4 indicate that (MMI) and trace-(MI) can simulate each other only by increasing the error probability $\alpha$, and that if $\mathbf{A}^{-1}$ is chosen to have a low effective rank, (MMI) can easily simulate trace-(MI) whereas trace-(MI) pays a price up to $d$ to simulate (MMI). Statements 5 and 6 imply respectively that, while one can always simulate quadratic form-(MI) with (MMI) by infinitesimally increasing the error probability $\alpha$, the converse is never true. Proposition 5.1 is proved in Section B.4.
Similar reasoning holds for (MVI) versus (VI) as well because the concentration regions, while holding time-uniformly or at a stopping time, remain in the same forms of $\mathcal{C}_{\mathrm{tr}}(\alpha)$, $\mathcal{C}_{\mathbf{v}}(\alpha)$, and $\mathcal{C}(\mathbf{A})$. We shall spell out the statistical consequences of these linear algebra results in Section 6.
5.2 Stepwise Scalarization
@@
As we mentioned earlier in Section 3.3, Howard et al. [2020, Section 6.5] prove, under different conditions, numerous inequalities for $\mathcal{S}_{d}$-valued martingale difference $\{\mathbf{Z}_{n}\}$ in the style of
@@
At this step, our methods in Section 3 and that of Howard et al. [2020] diverged. We have been able to directly construct the matrix-valued supermartingale
$\exp\left(-\frac{\mathbf{C}_{n}^{\prime}}{2}\right)\exp\left(\frac{\mathbf{Z}_{n}-\mathbf{C}_{n}}{2}\right)\dots\exp\left(\frac{\mathbf{Z}_{n}-\mathbf{C}_{n}}{2}\right)\exp\left(-\frac{\mathbf{C}_{n}^{\prime}}{2}\right)$ (68)
via Lemma 3.5, leading to a computable sequential test by our new (MVI). Howard et al. [2020], on the other hand, construct scalar e-processes from the condition (67) by applying a trace scalarization technique. Let us rephrase their “master lemma” behind this construction.
@@
The proof of this lemma can be found in Section B.5. We have numerous remarks regarding this lemma.
First, the upper bounded expression (70) is an instance of e-processes (if divided by $L_{0}=d$), introduced earlier in Section 3.4. The larger process (69), a supermartingale, can be used for a sequential test; whereas the smaller process (70), an e-process, leads to closed-form spectral time-uniform concentration inequalities (i.e. confidence sequences), but as a test process it is less powerful than (69).
Second, the divergence from the common assumption (67) to the matrix-valued (68) and scalar-valued (69) test supermartingale indicates that for some testing problems, two different routes can be pursued: our “matrix-native” approach, and the scalarization approach to the test process. Examples in Section 3.3 that fit in the assumption (67) include Examples 3.8 and 3.9 which Howard et al. [2020, Section 6.5] originally prove, but also Example 3.10: To see that Example 3.10 is in the form of (67),
$\exp(\gamma_{n}(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n}))=\exp(\underbrace{\gamma_{n}(\mathbf{X}_{n}-\mathbf{M}_{n})}_{:=\mathbf{Z}_{n}}+\underbrace{\gamma_{n}(\mathbf{M}_{n}-\widehat{\mathbf{X}}_{n})}_{:=\mathbf{C}_{n}}),$ (71)
which is exactly what is done by Wang and Ramdas [2024, Theorem 4.2] to derive a sharp closed-form matrix empirical Bernstein inequality. One of the crucial ways the two approaches differ is that with the scalar-valued test process (69), one rejects the null when the process surpasses $d/\alpha$; whereas with the matrix-valued test process (68), one rejects when the process $\npreceq\mathbf{A}$ where $\mathbf{A}$ is a matrix chosen freely beforehand, only subject to the “size constraint” that $\mathrm{tr}(\mathbf{A}^{-1})=\alpha$. Indeed, the extra degrees of freedom that come with $\mathbf{A}$ play a crucial role as we compare the two approaches empirically with the self-normalizing example (Example 3.8) in Section 6.
Third, the assumption (67) for scalarization being more restrictive than our Lemma 3.5, for some testing problems our concept of matrix-valued test supermartingales is unavoidable. These include Examples 3.7 and 3.12, to which the stepwise scalarization Lemma 5.2 is not applicable. Their one-step condition $\mathbb{E}(\mathbf{E}_{n}|\mathcal{F}_{n-1})\preceq\mathbf{F}_{n}$ leads naturally to a matrix-valued test supermartingale, and scalar-valued test supermartingales might be obtained by taking the trace or quadratic form after the formation of the matrix-valued one, which is not ideal according to Section 5.1.
5.3 Randomized Ahlswede-Winter Bounds
The uniformly randomized Ville’s inequality (8) by Ramdas and Manole [2024] is obtained by applying the uniformly randomized Markov’s inequality [Ramdas and Manole, 2024, Theorem 1.2]
$\mathbb{P}(X\leqslant Ua)\leqslant(\mathbb{E}X)a^{-1}$ (72)
where $U\sim\operatorname{Unif}_{(0,1)}$ independent from $X$. Here, we state a matrix version of (72), or a randomized version of the matrix Markov inequality (MMI).
Theorem 5.3 (Uniformly Randomized Matrix Markov’s Inequality).
Let $\mathbf{X}$ be a random matrix taking values in $\mathcal{X}\subseteq\mathcal{S}_{d}^{+}$, and $U\sim\operatorname{Unif}_{(0,1)}$ independent from $\mathbf{X}$. For any $\mathbf{A}\in\mathcal{S}_{d}^{++}$,
$\mathbb{P}(\mathbf{X}\npreceq U\mathbf{A})\leqslant\mathrm{tr}((\mathbb{E}\mathbf{X})\mathbf{A}^{-1}).$ (73)
It is clear that (73) generalizes the scalar (72) and tightens (MMI). The randomization factor $U$, we note, can be any random variable that is stochastically larger than $\operatorname{Unif}_{(0,1)}$ (“super-uniform”), that is, $\mathbb{P}(U\leqslant x)\leqslant x$ for all $x\geqslant 0$. More generally, we shall present and prove a stronger version of Theorem 5.3 in Appendix A, where the scalar randomization factor $U$ is further replaced by a PSD random matrix whose distribution belongs to a class of “super-uniform” measures on $\mathcal{S}_{d}^{+}$. The generalized version of Theorem 5.3 in Appendix A also reveals the equality condition for (MMI), randomized or not.
Theorem 5.3 implies that various matrix concentration inequalities by Ahlswede and Winter [2002] can be tightened via a randomization factor as well. For example, the Chebyshev inequality (10) for $n$ i.i.d. matrices can be randomized as follows.
Corollary 5.4 (Uniformly Randomized Matrix Chebyshev’s Inequality).
Let $\mathbf{X}_{1},\dots,\mathbf{X}_{n}$ be random matrices taking values in $\mathcal{S}_{d}$ with common mean matrix $\mathbb{E}\mathbf{X}_{i}=\mathbf{M}$ and variance matrix upper bound $\mathbf{Var}(\mathbf{X}_{i})\preceq\mathbf{V}$, such that $\mathbb{E}\{\mathbf{X}_{i}-\mathbf{M},\mathbf{X}_{j}-\mathbf{M}\}=0$ for any $i\neq j$; $U$ a super-uniform random scalar independent from them all. Then, letting $\overline{\mathbf{X}}_{n}=\frac{1}{n}(\mathbf{X}_{1}+\dots+\mathbf{X}_{n})$, for any $\mathbf{A}\in\mathcal{S}_{d}^{++}$.
$\mathbb{P}(\operatorname{abs}(\overline{\mathbf{X}}_{n}-\mathbf{M})\npreceq\sqrt{U}\mathbf{A})\leqslant n^{-1}\mathrm{tr}(\mathbf{V}\mathbf{A}^{-2}),$ (74)
We omit the proof since it is a straightforward application of Theorem 5.3 to the moment bound Lemma B.2. We next discuss the implication for (MVI) as the randomization here raises some subtleties in its two equivalent forms, the time-uniform one and the stopped one.
Theorem 5.5 (Uniformly Randomized Matrix Ville’s Inequality).